OpenStudy (anonymous):

what is the inverse of y=x^2? and is it a function? what restrictions are placed in the inverse of y=x^2 to make it an inverse?

OpenStudy (anonymous):

HELP PLEASE!

OpenStudy (accessdenied):

If you switch the y and x and solve for x... y = x^2 <=> x = y^2 sqrt(x) = |y| +- sqrt(x) = y This is not one, but two functions.

OpenStudy (accessdenied):

solve for y, rather

OpenStudy (anonymous):

Thankyou for your help :) I understand it now!

OpenStudy (accessdenied):

No problem! You also understand the restriction part?

OpenStudy (anonymous):

not really :/

OpenStudy (accessdenied):

Well, if we restricted the domain (our x-values) of our original function to be only positive real numbers and 0, we'd get this half of the parabola:  Since we no longer have the other half of the graph, this satisfies the horizontal line test.